OpenStudy (anonymous):

when a particle is located a distance x meters from the origin a force of cos(pi*x/5) newtons acts on it. find the work done in moving the particle from x=2 to x=2.5

OpenStudy (anonymous):

\[ \Large W = \int_{x_i}^{x_f}F(x)dx = \int_2^{2.5}\cos\left(\frac{\pi x}{5}\right)dx \]

OpenStudy (anonymous):

thanks!